OpenStudy (anonymous):

For what value of b will the polynomial p(x)= -2x^3+bx^2-5x+2 have the same remainder when it is divided by x-2 and x+1

hartnn (hartnn):

to get the remainder when p(x) is divided by x-2, just find p(2) know how to ? similarly, to get the remainder when p(x) is divided by x+1, just find p(-1) and since they are equal, you equate p(2) = p(-1) got this ?

OpenStudy (anonymous):

oh omg okay ill try this

OpenStudy (anonymous):

I got 5, but the answr is 11

hartnn (hartnn):

The Remainder Theorem: When you divide a polynomial f(x) by x-c the remainder r will be f(c)

hartnn (hartnn):

ok let me try

OpenStudy (anonymous):

Okay

hartnn (hartnn):

when x=-1, p(-1) = b+9

hartnn (hartnn):

when x=2, p(2) = 4b-24

hartnn (hartnn):

so, 4b-24 = b+9 3b = 9+24 b=33/3

hartnn (hartnn):

where did you made an error ?

OpenStudy (anonymous):

I somehow got 3b=-9+24

OpenStudy (anonymous):

Maybe mis calculated. but I understand how you did it. Thank you so much

hartnn (hartnn):

so your when x=-1, p(-1) = b-9 instead of b+9 ?

OpenStudy (anonymous):

Actually -b+9. I have to re-do that part, I think I messed up by accident somewhere

OpenStudy (anonymous):

Ah. b(-1)^2 should have been positive b but I wrote negative by mistake

hartnn (hartnn):

ok ok, glad that you got it :) welcome ^_^

OpenStudy (anonymous):

:)